# Halfway to 3D: Ensembling 2.5D and 3D Models for Robust COVID-19 CT Diagnosis

### 16 Mar 2026 v1

@@ Abstract
We propose a deep learning framework for COVID-19 detection and disease classification from chest CT scans that integrates both 2.5D and 3D representations to capture complementary slice-level and volumetric information. The 2.5D branch processes multi-view CT slices (axial, coronal, sagittal) using a DINOv3 vision transformer to extract robust visual features, while the 3D branch employs a ResNet-18 architecture to model volumetric context and is pretrained with Variance Risk Extrapolation (VREx) followed by supervised contrastive learning to improve cross-source robustness. Predictions from both branches are combined through logit-level ensemble inference. Experiments on the PHAROS-AIF-MIH benchmark demonstrate the effectiveness of the proposed approach: for binary COVID-19 detection, the ensemble achieves 94.48% accuracy and a 0.9426 Macro F1-score, outperforming both individual models, while for multi-class disease classification the 2.5D DINOv3 model achieves the best performance with 79.35% accuracy and a 0.7497 Macro F1-score. These results highlight the benefit of combining pretrained slice-based representations with volumetric modeling for robust multi-source medical imaging analysis. Code is available at https://github.com/HySonLab/PHAROS-AIF-MIH.
@@ 4.5 Results
We evaluate the 2.5D multi-view model based on the DINOv3 backbone on the binary COVID-19 detection task. The model achieves a validation accuracy of 91.23% and a mean Macro F1-score of 0.7986, demonstrating strong performance using multi-view slice representations. Table 5 shows the per-source F1-score across the four data sources.
@@
For the multi-class classification task (A, G, COVID, Normal), the 2.5D DINOv3-based model achieves an overall validation accuracy of 79.35% and a mean Macro F1-score of 0.7422. Table 6 presents the class-wise performance of the model.
To assess fairness across demographic groups, we analyze performance across gender. The model achieves an F1-score of 0.8562 for male and 0.6283 for female patients. The performance difference suggests the presence of domain shifts between institutions, which remains a challenge for multi-source medical imaging models.
@@ 4.5.3 Ensemble Results
For the multi-class disease classification task, the ensemble achieves an overall validation accuracy of 78.06% and a mean Macro F1-score of 0.7321. While the ensemble improves stability across sources, the best overall performance for this task is still achieved by the 2.5D DINOv3 model.
@@
A 0.8099 0.8108 0.7928
G 0.2069 0.4737 0.4211
COVID 0.8354 0.8608 0.8608
Normal 0.8395 0.8537 0.8537
@@ 4.6 Discussion
While the ensemble demonstrates robust generalization across most data sources, we observe a severe performance degradation on Source 2, where the 3D model’s Macro F1-score falls to 0.4828. This sharp decline is primarily attributable to pronounced domain shift and class imbalance unique to this subset. Unlike other sources, Source 2 likely features distinct scanner acquisition protocols or varying slice thicknesses that disrupt the volumetric spatial continuity upon which the 3D ResNet-18 relies. Furthermore, the inherent morphological difficulty in distinguishing Squamous Cell Carcinoma (Class G) from other pathologies severely exacerbates these errors. This suggests that while global 3D volumetric features capture excellent broader context, they are significantly less resilient to severe spatial distribution shifts than 2.5D slice-level representations.

### 18 Mar 2026 v2

@@ Abstract
We propose a deep learning framework for COVID-19 detection and disease classification from chest CT scans that integrates both 2.5D and 3D representations to capture complementary slice-level and volumetric information. The 2.5D branch processes multi-view CT slices (axial, coronal, sagittal) using a DINOv3 vision transformer, while the 3D branch employs a ResNet-18 architecture pretrained with Variance Risk Extrapolation (VREx) and further refined with supervised contrastive learning to improve cross-source robustness. Predictions from both branches are combined via logit-level ensembling.
Experiments on the PHAROS-AIF-MIH benchmark demonstrate the effectiveness of the proposed approach. On the test set, our method achieved runner-up in the Multi-Source COVID-19 Detection Challenge, with the best ensemble reaching a Macro F1-score of 0.751. For the Fair Disease Diagnosis Challenge, our approach ranked third place, achieving a best Macro F1-score of 0.633 with improved performance balance across genders. These results highlight the benefits of combining pretrained slice-based representations with volumetric modeling, as well as the importance of ensemble strategies for improving robustness and fairness in multi-source medical imaging tasks. Code is available at https://github.com/HySonLab/PHAROS-AIF-MIH.
@@ 4.5 Results
We evaluate the 2.5D multi-view model based on the DINOv3 backbone on the binary COVID-19 detection task. The model achieves a validation accuracy of 93.51% and a mean Macro F1-score of 0.8221, demonstrating strong performance using multi-view slice representations. Table 5 shows the per-source F1-score across the four data sources.
@@
For the multi-class classification task (A, G, COVID, Normal), the 2.5D DINOv3-based model achieves an overall validation accuracy of 76.77% and a mean Macro F1-score of 0.7229. Table 6 presents the class-wise performance of the model.
To assess fairness across demographic groups, we analyze performance across gender. The model achieves an F1-score of 0.7848 for male and 0.6611 for female patients. The performance difference suggests the presence of domain shifts between institutions, which remains a challenge for multi-source medical imaging models.
@@ 4.5.3 Ensemble Results
For the multi-class disease classification task, the ensemble achieves an overall validation accuracy of 76.77% and a mean Macro F1-score of 0.7229. While the ensemble improves stability across sources, the best overall performance for this task is still achieved by the 2.5D DINOv3 model.
@@
A 0.8099 0.7778 0.7778
G 0.2069 0.4390 0.4390
COVID 0.8354 0.8537 0.8537
Normal 0.8395 0.8537 0.8354
4.6 Test Submission Results
Our methods were evaluated on the official test server of the PHAROS-AIF-MIH benchmark. The proposed approach achieved strong rankings across both tasks, demonstrating robustness to multi-source distribution shifts and fairness considerations.
Multi-Source COVID-19 Detection (Task 1).
Our team achieved runner-up position in the challenge. Table 11 summarizes the performance across different hospitals. The 2.5D model outperforms the standalone 3D model, while ensembling further improves robustness. The best performance is obtained with the 0.5-weighted ensemble, achieving a Macro F1-score of 0.751.
Table 11: Test results for Task 1 across hospitals.
Method Avg H1 H2 H3 H4
2.5D 0.741 0.910 0.691 0.493 0.869
3D 0.699 0.825 0.608 0.495 0.868
Ens (0.3) 0.739 0.895 0.673 0.496 0.892
Ens (0.5) 0.751 0.914 0.712 0.495 0.883
Ens (0.7) 0.744 0.912 0.696 0.495 0.873
The results show that ensembling improves generalization across most hospitals, particularly for H2, while performance on H3 remains challenging for all methods.
Fair Disease Diagnosis (Task 2).
Our team ranked third place in this task. Table 12 reports the Macro F1-scores along with gender-specific performance. The best result is achieved by the 0.7-weighted ensemble, with a Macro F1-score of 0.633.
Table 12: Test results for Task 2 with gender-based evaluation.
Method Avg Female Male
2.5D 0.555 0.709 0.402
3D 0.572 0.756 0.388
Ens (0.3) 0.568 0.732 0.403
Ens (0.5) 0.561 0.718 0.403
Ens (0.7) 0.633 0.709 0.557
The results highlight a noticeable performance gap between genders, with higher performance on female samples for most methods. However, the 0.7-weighted ensemble significantly improves male performance, leading to the best overall Macro F1-score. This suggests that adjusting ensemble weights can help balance fairness and overall accuracy.
@@ 4.7 Discussion
While the ensemble demonstrates robust generalization across most data sources, we observe a severe performance degradation in distinguishing Squamous Cell Carcinoma (Class G) from other pathologies severely exacerbates these errors. This suggests that while global 3D volumetric features capture excellent broader context, they are significantly less resilient to severe spatial distribution shifts than 2.5D slice-level representations.